OpenStudy (anonymous):

Given: a = 8, b = 6, Angle C = 60 Angle A = 14° 46° 74° 106°

OpenStudy (phi):

we could use the law of cosines to find side c, and then the law of sines to find angle A

OpenStudy (anonymous):

i The only angle you are given is C, so use c^2 = a^2 + b^2 - 2abcosC...now substitute

OpenStudy (anonymous):

Would you use the formula a^2=b^2+c^2-2bcCosA

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay thanks ill try

OpenStudy (anonymous):

okay

OpenStudy (phi):

This is a 2 step problem. find side c using \[ c^2 = a^2 + b^2 - 2\ a\ b \cos(C) \] after you find c, you can use the Law of Sines \[ \frac{\sin A}{a}= \frac{\sin C}{c} \]

OpenStudy (anonymous):

okay thanks